OpenStudy (anonymous):

How to solve x ^ -3/2 = 1/729

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

\[x^{-3/2} = \frac{1}{729}\]First you want to change \(x^{-3/2}\) to havign a positive exponent, and you do hat by putting 1 over it.

OpenStudy (jhannybean):

\[\frac{1}{x^{3/2}}= \frac{1}{729}\]

OpenStudy (jhannybean):

Now change \(x^{3/2}\) changing it to square root form \(\implies \sqrt{x^3}\) since the base of any square root is always, always 2 unless otherwise stated.

OpenStudy (jhannybean):

\[\frac{1}{\sqrt{x^3}} = \frac{1}{729}\] To get rid of our square root we must square the entire fraction. Squaring a square root "undoes" the square root in a sense.\[\left(\frac{1}{\sqrt{x^3}}\right)^2 = \left(\frac{1}{729}\right)^2 \implies \frac{1^2}{(\sqrt{x^3})^2} = \frac{1^2}{729^2}\]

OpenStudy (jhannybean):

Are you following? @Stephanie1000000